On the surface of the ice s = 20.4m for sliding stones at initial velocity v = 3m / s
stopped on the road Find the coefficient of friction of stones on ice
1
Expert's answer
2020-03-09T11:17:56-0400

The change of kinetic energy is equal to work done by a friction force


"\frac{mv^2}{2}=\mu mgs."

Hence


"\mu=\frac{v^2}{2gs}=\frac{3^2}{2\times 9.8\times 20.4}=0.023"